OpenStudy (anonymous):

Kyle ran 2/3 miles on Monday, 1 1/6 miles on Tuesday, and 7/12 miles on Wednesday. How many total miles did Kyle run? Express your answer in simplest form. @ttb123456789 @iGreen @PaulaLovesSchool13 @TexasBoy123 @Rainbow_dashie please help fan and medal

OpenStudy (igreen):

They all have a common denominator of 12. \(\dfrac{2}{3} + \dfrac{7}{6} + \dfrac{7}{12} \rightarrow \dfrac{8}{12} + \dfrac{14}{12} + \dfrac{7}{12} \rightarrow \dfrac{8+14+7}{12}\)

OpenStudy (anonymous):

29?

OpenStudy (anonymous):

@iGreen am I right?

OpenStudy (igreen):

Yes, that gives us \(\dfrac{29}{12}\). Does it want you to put it in mixed number form?

OpenStudy (anonymous):

yes

OpenStudy (igreen):

Can you do that?

OpenStudy (anonymous):

no

OpenStudy (igreen):

First you divide 29 by 12, which gives you 2 and a remainder of 5. So the whole number will be 2, and the numerator will be 5, and we keep the denominator of 12. So what does that give us?

OpenStudy (anonymous):

2 5/12?